### v3

# Glauber predictions for oxygen and neon collisions at energies available at the LHC

Abstract
The Glauber model is a widely used framework for describing the initial conditions in high-energy nuclear collisions. TGlauberMC is a Monte Carlo implementation of this model that enables detailed, event-by-event calculations across various collision systems. In this work, I present an updated version of TGlauberMC (v3.3), which incorporates recent theoretical developments and improved parameterizations, especially relevant for small collision systems. I focus on the oxygen–oxygen (OO), neon–neon (NeNe), and proton-oxygen (pO) collisions at the Large Hadron Collider (LHC) in July 2025, where precise modeling of nuclear geometry and fluctuations is essential. The updated version includes revised nuclear density profiles and an enhanced treatment of nucleon substructure. Geometrical cross sections for all relevant collision systems are calculated and initial-state observables are explored to provide predictions for particle production trends at $\sqrt{s_{\rm NN}}=5.36$ TeV. In particular, a prediction for the centrality dependence of mid-rapidity multiplicity in OO and NeNe collisions is obtained. The updated code is publicly available to support the heavy-ion community with a robust and flexible tool for studying strongly interacting matter in small and intermediate-sized nuclear systems.
Revisions && changes of the arXiv document and code:
v1
, 08 Jul 2025: initial document, code v3.3
v2, 03 Sep 2025: submitted version, code v3.3.1 includes fix for Opar2 and new NNLOsat profile (“Osat”)
v3, 03 Feb 2026: published version, code v3.3.2 includes smearing with TRENTO, and static functions
I Introduction
The study of heavy-ion collisions at ultra-relativistic energies has been instrumental in exploring the properties of the quark-gluon plasma (QGP), a state of matter believed to have existed shortly after the Big Bang Busza et al. (2018). Experiments at the Relativistic Heavy Ion Collider (RHIC) and the LHC have provided compelling evidence for the formation of QGP in large collision systems such as gold–gold (AuAu) and lead–lead (PbPb) at high collision energies Arsene and others (2005); Adcox and others (2005); Back and others (2005); Adams and others (2005); Acharya and others (2024b); Hayrapetyan and others (2025c). Among the most prominent observables signaling QGP formation are collective flow phenomena, characterized by anisotropic flow coefficients such as $v_{2}$ and higher moments Heinz and Snellings (2013), and the phenomenon of jet quenching, in which high-energy partons lose energy as they traverse the dense medium Cunqueiro and Sickles (2022). Surprisingly, several QGP-like features have also been observed in small collision systems, such as proton-lead (pPb) and even high-multiplicity proton-proton (pp) collisions Loizides (2016a); Nagle and Zajc (2018); Grosse-Oetringhaus and Wiedemann (2024). These include long-range correlations and non-zero elliptic flow coefficients, raising questions about the minimal conditions required for QGP formation. However, one signature remains notably absent in small systems: jet quenching. Despite observing various collective phenomena, experiments have not identified energy loss of high-$p_{T}$ partons in pPb collisions Acharya and others (2018); Aad and others (2023). To further elucidate the system-size dependence of QGP signatures and to search for the possible onset of jet quenching in intermediate-sized systems, oxygen-oxygen (OO) and neon-neon (NeNe) collisions at a center-of-mass energy ($\sqrt{s_{\rm NN}}$) of 5.36 TeV per nucleon pair provided by the LHC in July, 2025 are of great interest Citron and others (2019); Brewer et al. (2021). These collision systems provide a crucial bridge between small and large systems, allowing precise control measurements of observables such as jet quenching and elliptic flow Lim et al. (2019); Sievert and Noronha-Hostler (2019); Huang et al. (2020); Huss et al. (2021); Jia and others (2024). Specifically, the role of $\alpha$ clusters in small nuclei including 16O has long been discussed and may be imprinted in the final-state flow measurements Rybczyński et al. (2018); Rybczyński and Broniowski (2019); Li et al. (2020); Summerfield et al. (2021); Ding et al. (2023); Wang et al. (2024); Zhang et al. (2025). Recently, comparing $v_{2}$ in NeNe to $v_{2}$ in OO collisions was highlighted to provide a precision test of the role of initial geometry and the QGP paradigm Giacalone and others (2025). Additionally, proton–oxygen (pO) collisions at a $\sqrt{s_{\rm NN}}=9.62$ TeV were provided, primarily to improve hadronic interaction models used in the interpretation of extensive air showers in cosmic-ray physicsDembinski et al. (2020), but they also present a unique opportunity to investigate QGP-related phenomena.
Interpretation of high-energy nucleus collision data typically depends on the use of classical Glauber Monte Carlo (MCG) models to describe the initial condition Miller et al. (2007); d’Enterria and Loizides (2021). The general approach is to deduce properties of QCD matter by measuring the distributions of a variety of observables in AA collisions, and comparing them to the same measurements in pp collisions, where a QGP is not expected to be produced. To compare colliding systems of different sizes requires hence appropriate normalization of the measured distributions by using e.g. the number of participating nucleons ($N_{\rm part}$), the number of independent binary nucleon-nucleon (NN) collisions ($N_{\rm coll}$), the medium transverse area ($A_{\perp}$), or the overlap eccentricity $\epsilon_{n}$ (given by the $n^{\rm th}$ moments of its azimuthal spatial distribution).
Widely used and publicly available MCG models are the TGlauberMC Loizides et al. (2015); Loizides (2016b); Loizides et al. (2018), which developed from the original PHOBOS implementation Alver et al. (2008) and GLISSANDO codes Broniowski et al. (2009); Rybczynski et al. (2014); Bożek et al. (2019). The Glauber initial condition is widely used in simulations of heavy-ion collisions as provided by TRENTO Moreland et al. (2015), Trajectum Nijs and van der Schee (2022b) and MUSIC Paquet et al. (2016), as well as by event generators such as HIJING Wang and Gyulassy (1991), AMPT Lin et al. (2005), URQMD Bass and others (1998) and EPOS Werner et al. (2010). An alternative description of the initial condition is provided by IP-Glasma Schenke et al. (2019).
In this article, an updated version (v3.3) of the TGlauberMC model is used to predict properties of OO, NeNe, as well as pPb and PbPb collisions at 5.36 TeV, and pO collisions at 9.62 TeV. In particular, I focus on the calculation of the expected cross sections and discuss aforementioned associated initial-state quantities, such as $N_{\rm part}$, $N_{\rm coll}$ and eccentricity distributions for OO and NeNe collisions including the ratio of eccentricities for NeNe to OO as a proxy for the respective ratio of $v_{2}$ in these two systems. Finally, particle production at 5.36 TeV is discussed by outlining some features derived from the multiple parton interaction (MPI) picture Sjöstrand (2018), and use existing data at 5.36 TeV to predict the multiplicity in OO and NeNe collisions. As for previous versions of the model, the source code for TGlauberMC (v3.3) has been made publicly available at HepForge 92.
The remainder of the article is divided in the following sections. Section II briefly introduces the general MCG approach. Section III focuses on the treatment of the transverse nucleon–nucleon overlap. Section IV discusses the available parameterizations and calculations of 16O and 20Ne nuclear densities. Section V presents the predictions for the hadronic cross sections at 5.36 and 9.62 TeV. Section VI discusses the $N_{\rm part}$, $N_{\rm coll}$ and eccentricity distributions for OO and NeNe. Section VII focuses on particle production using the MPI model.. Section VIII provides a concise summary of the main findings of this article. Appendix A describes the improvements of TGlauberMC (v3.3) and provides a short user’s guide. Appendix B introduces a new software library called “Trajectum Nucleus Generator (TrNucGen)” to provide nuclear densities calculated by Trajectum. Appendix C provides a set of extra figures not available in the published version.
II The Monte Carlo Glauber approach
Monte Carlo Glauber (MCG) calculations evaluate the cross section of two nuclei with $A$ and $B$ nucleons stochastically by distributing them in coordinate space according to the corresponding nuclear densities, separated by an impact parameter $b$ sampled from $\mathrm{d}\sigma/\mathrm{d}b\propto b$. For simplicity, the impact-parameter variable $b$ or $b_{\rm NN}$ denotes the distance between nuclei (or nucleons) in the transverse plane. The nuclear transverse profiles of the nuclei are usually approximated with parametrizations of charge density distributions extracted from low-energy electron-nucleus scattering experiments De Vries et al. (1987). For large nuclei, they are usually described by a 3-parameter Fermi (3pF) distribution
$\rho(r)=\rho_{0}\left(1+w\frac{r^{2}}{R^{2}}\right)/\left(1+\exp(\frac{r-R}{a})\right)$ (1)
with half-density radius $R$ and diffusivity $a$, and $w\neq 0$ describing small deviations from a spherical shape. Following the eikonal approximation Capel (2020), the collision is treated as a sequence of independent binary nucleon-nucleon collisions, i.e. the nucleons travel on straight-line trajectories and their interaction probability does not depend on the number of collisions they have suffered before. In the so-called “hard-sphere” (HS) approximation, an interaction takes place between two nucleons if the transverse distance $d$ (or generally the transverse nucleon–nucleon impact parameter $b_{\rm NN}$) between their centers satisfies $d<\sqrt{\sigma_{\rm NN}/\pi}$.
In the MCG approach, the number of participants and collisions are computed by counting the number of nucleons with at least one collision, and the total number of nucleon–nucleon collisions, respectively. The collision cross section is the fraction of simulated collisions with $N_{\rm coll}>0$ of all simulated collisions. Quantities such as the eccentricity Alver and others (2007), the triangularity Alver and Roland (2010), and higher moments Teaney and Yan (2011) of the collision region at impact parameter $b$ are given by
$\varepsilon_{n}(b)=\frac{\left<r^{n}\cos(n\phi-n\psi)\right>}{\left<r^{n}\right>}$ (2)
where $n$ denotes the moment, $r=\sqrt{x^{2}+y^{2}}$ and $\psi=\tan^{-1}\frac{y}{x}$. The averages are computed by considering the central positions of either, participant nucleons or binary nucleon–nucleon collisions, or of an admixture of the two, and additionally may be smeared to model matter production Alver and others (2008).
Figure 1: Nucleon–nucleon impact parameter distribution (left) and interaction probability (right) at 5.36 TeV shown for the different transverse overlap functions ($\Gamma$-parameterization with $0\leq\omega\leq 1$, TRENTO with $0.4\leq w\leq 0.7$, HIJING and PYTHIA tune Monash) as discussed in the text. The thin dashed line indicates the hard-sphere approximation.
III Transverse overlap function
Several extensions to the hard-sphere approximation have been explored to account for the fact that the nucleon matter profile is not a black disc d’Enterria et al. (2010). A commonly applied extension, referred to as TRENTO Moreland et al. (2015), is to assume the nuclear matter profile of a nucleon to be a Gaussian with a certain width $w$, so that the nuclear overlap function
$T_{\rm NN}(b_{\rm NN})\propto\exp\left[-\left(b_{\rm NN}/B\right)^{m}\right]$ (3)
is also Gaussian, i.e. $m=2$ and $B=2w$. For some settings (“bProfile = 3” Sjöstrand (2018)) PYTHIA assumes $B=1$ fm as characteristic size of the proton, and uses $m$ to interpolate between Exponential ($m\to 1$) and Gaussian ($m\to 2$) shape. The widely-used “Monash” tune sets $m=1.85$ Skands et al. (2014). HIJING Wang and Gyulassy (1991) uses an overlap function based on modified Bessel functions of the 3rd kind,
$T_{\rm NN}({b_{\rm NN}})\propto(\mu\,b_{\rm NN})^{3}\,K_{3}(\mu\,b_{\rm NN})\,$, (4)
with $\mu={1.5}/{\sqrt{\sigma_{\rm NN}}}$. In the Eikonal formalism, the probability to interact is then given as $p(b_{\rm NN})=1-\exp\left(-k\,T_{\rm NN}(b_{\rm NN})\right)$ where $k$ is obtained by requiring $2\pi\,\int p(b_{\rm NN})\,b_{\rm NN}\,{\rm d}b_{\rm NN}=\sigma_{\rm NN}$. At LHC collision energies, a practical parametrization Rybczyński and Wlodarczyk (2013) based on the Euler $\Gamma(z)$ and incomplete $\Gamma(\alpha,z)$ functions, in the following also called $\Gamma$-parametrization, is
$p(b_{\rm NN})\propto\Gamma\left(1/\omega,\frac{b_{\rm NN}^{2}}{d^{2}\omega}\right)\bigg/\Gamma(1/\omega)\,$, (5)
where $\omega$ interpolates between the hard-sphere ($\omega\to 0$) and Gaussian ($\omega\to 1$) cases.
The different choices for the nucleon–nucleon impact parameter distributions and corresponding interaction probabilities are shown in Fig. 1 for the $\Gamma$-parameterization with $0\leq\omega\leq 1$, TRENTO with $0.4\leq w\leq 0.7$, HIJING and PYTHIA tune Monash. Beyond the hard-sphere approximation, there is a quite a large parameter space, leading to a significant uncertainty resulting from the dependence on the parameterization. However, HIJING, PYTHIA, as well as the $\Gamma$-parameterizations for $\omega\sim 0.5$ and TRENTO for $w\sim 0.5$ lead to relatively similar distributions, quite distinct from the hard-sphere approximation.
Figure 2: Nuclear charge density for 16O versus radius for various descriptions compared to data from electron scattering Sick and Mccarthy (1970). As explained in the text, “Oho2” is a parameterization based on the HO model, whose parameters are fit to the data, while “Odat” was already provided in Sick and Mccarthy (1970). “Osat” is the exact result of the NNLOsat calculation Somà et al. (2020)), and “Opar2” the respective 3pF fit. “Opar” and “Oho” profiles were previously included in TGlauberMC. The distributions are normalized as $\int\rho\,r^{2}\,{\rm d}r=Ze$. The respective $r_{\rm rms}$ values are given in the legend, denoted as $s$.
IV Nuclear density profiles for oxygen and neon
The standard method to construct the nucleon configurations for MCG calculations is to sample the nucleon positions from $r^{2}\rho$, where the nuclear density $\rho$ may include angular deformations Loizides et al. (2015). To mimic a hard-core repulsion potential between nucleons, a minimum inter-nucleon separation ($d_{\rm min}$) of $0.4$ fm between their centers is enforced when sampling the positions of the nucleons inside a nucleus. In order to ensure that the center-of-mass of each constructed nucleus is at $(0,0,0)$, the nucleons are individually “shifted” by the necessary amount, after all nucleon positions were sampled. Both, imposing the hard-core repulsion as well as the re-centering, induce a bias from the original $\rho$ distribution. As discussed in Ref. Loizides et al. (2018), this can be overcome by a “reweighting” procedure, modifying the original $\rho$ distribution to counteract the bias. However, for $d_{\rm min}=0.4$ fm, and recentering via rotation around the $z$-axis (recentering option 4 in TGlauberMC), residual differences are negligible.
Table 1: Nuclear parameters for oxygen and neon parameterizations provided by TGlauberMC (v3.3) with given names as in the code. The 3pF parameterization is given in Eq. (1). The HO parameterization is given in Eq. (6). The last two configurations used deformed profiles, with deformation values as given in the comment field, and defined in Trajectum Giacalone and others (2025). The prefix “TR” indicates the profile is taken from Trajectum, where the name following the prefix denotes the system and version number as used in Trajectum. Included as “Osat” is also the exact profile from the ${\rm NNLO}_{\rm sat}$ chiral Hamiltonian calculation Somà et al. (2020). For completeness, “O” is also listed which uses explicit nucleon configurations Lonardoni et al. (2018) as introduced in Lim et al. (2019).
Name AA $Z$ $R$ [fm] or $\alpha$ $a$ [fm] $w$ Comments
Odat 16 8 2.608 0.513 -0.051 Original parametrization (3pF plus corrections) from Sick and Mccarthy (1970)
Opar 16 8 2.608 0.513 -0.051 3pF parameterization De Vries et al. (1987) (same as TR_OXYGEN_V3)
Opar2 16 8 1.850 0.497 0.912 3pF parameterization (TR_OXYGEN_V4, fit to ${\rm NNLO}_{\rm sat}$ Somà et al. (2020))
Oho 16 8 1.544 1.833 - Harmonic Oscillator De Vries et al. (1987)
Oho2 16 8 1.506 1.819 - Harmonic Oscillator (new fit to data from Sick and Mccarthy (1970))
Osat 16 8 - - - Exact NNLOsat profile from Somà et al. (2020), use instead of Opar2
O 16 8 - - - Nucleon configurations Lim et al. (2019) (same as in TGlauberMC v2.7)
Ne 20 10 2.805 0.571 - 2pF parametrization De Vries et al. (1987)
Ne2 20 10 2.740 0.572 - 2pF parametrization De Vries et al. (1987)
Ne3 20 10 2.791 0.698 -0.168 3pF parametrization De Vries et al. (1987)
NeTr2 20 10 2.805 0.571 - TR_NEON_V2, (same as Ne with $\beta_{2}=0.721$ from Pritychenko et al. (2016))
NeTr3 20 10 2.724 0.498 - TR_NEON_V3, $\beta_{2}=0.4899$, $\beta_{3}=0.2160$, $\beta_{4}=0.3055$
Figure 2 shows the nuclear charge density for 16O for data from electron scattering Sick and Mccarthy (1970), including the original parameterization (“Odat”), as well as 3pF distributions with various parameters available in the literature (“Opar”, and “Opar2”). Ref. De Vries et al. (1987) also gives a parameterization based on the (modified) Harmonic Oscillator (HO) model
$\rho(r)=\rho_{0}\left(1+\alpha\left(x/a\right)^{2}\exp{-\left(x/a\right)^{2}}\right)$ (6)
labeled “Oho” in the figure. For this work, I also provide a new HO fit to the data, called “Oho2”, leading to $\alpha=1.506\pm 0.023$ and $1.819\pm 0.004$ with $\chi^{2}/n_{\rm dof}\approx 1$ assuming about 1% relative error on the data points. While it overall has a worse $\chi^{2}$ than “Odat”, it only has two parameters and can be used as alternative and allows systematic variation using the uncertainties provided on the fit parameters. To reflect the charge density, the distributions shown in Fig. 2 are normalized as $\int\rho\,r^{2}\,{\rm d}r=Ze$. A detailed list of all parameterizations shown in Fig. 2 is given in Tab. 1. The $r_{\rm rms}=\sqrt{\langle r\rangle}$ values are reported in the legend of the figure, and differ by nearly up to 5%. The reason is that the data do not constrain the tail of the distribution, which was found to affect the $r_{\rm rms}$ by a few percent, and thus experimental results between $2.65\pm 0.04$ and $2.73\pm 0.03$ have been reported Sick and Mccarthy (1970). Furthermore small differences can also be expected from inclusion of neutrons in the density distributions, and the use of point-like protons ignoring the proton form factor. However, the distribution “Opar2” which results from a fit to the ${\rm NNLO}_{\rm sat}$ calculation Somà et al. (2020) exhibits a significantly larger $r_{rms}$ than the data (and than reported in Somà et al. (2020)), and hence is not further considered here. Instead, the exact nuclear density obtained from the ${\rm NNLO}_{\rm sat}$ chiral Hamiltonian calculation Somà et al. (2020) is included and also shown in the figure. The nuclear charge density directly calculated by ${\rm NNLO}_{\rm sat}$ perfectly agrees with the data (also shown in the figure).
Table 2: Oxygen-16 and Neon-20 NLEFT and PGCM calculations taken from Trajectum Giacalone and others (2025).
Profile Description
TR_OXYGEN_V12 NLEFT pinhole, $\pm$ weights, with periodicity ambiguities resolved
TR_OXYGEN_V14 PGCM, constraints imposed on projected states, without explicit $\alpha$ clustering ($C=0$)
TR_OXYGEN_V15 PGCM, constraints imposed on projected states, including explicit $\alpha$ clustering ($C=1$)
TR_NEON_V11 NLEFT pinhole, $\pm$ weights, with periodicity ambiguities resolved
TR_NEON_V13 PGCM, constraints imposed on projected states, without explicit $\alpha$ clustering ($C=0$)
TR_NEON_V14 PGCM, constraints imposed on projected states, with explicit $\alpha$ clustering ($C=1$)
Instead of sampling nucleon positions from average $\rho$ distributions, explicit calculations of the nucleon configuration based on the many-body wave function, is desired to account for inter-nucleon correlations, in particular important for small and deformed nuclei Dobaczewski et al. (2025). In this case, an explicit $d_{\rm min}$ criterion between nucleons should not be imposed. Previous work Lim et al. (2019) used an ab-initio quantum Monte Carlo calculation of the Oxygen nucleon configurations with two- and three-body local chiral potentials at next-to-next-to-leading order Lonardoni et al. (2018), which were available in TGlauberMC (since v2.7, named “O”, see Tab. 1). Recently, Oxygen and Neon configurations where derived within the framework of Nuclear Lattice Effective Field Theory (NLEFT) and the ab-initio Projected Generator Coordinate Method (PGCM), and used as input for the initial conditions in Trajectum Giacalone and others (2025). The NLEFT framework is well suited to probe collective phenomena in the ground states of nuclei Lu et al. (2019). Similarly to Summerfield et al. (2021), the positions of the nucleons were traced using the pinhole Elhatisari et al. (2017) algorithm. In this way, the NLEFT configurations, which can be obtained keeping track of positive and negative weights, and with resolving ambiguities introduced by the underlying lattice periodicity, capture short-range correlations without explicitly enforcing $\alpha$-clusters. In the PGCM approach Frosini et al. (2022); Hüther et al. (2020), a set of mean-field states with different quadrupole and octupole deformations is first generated to explore a range of possible nuclear shapes. Each of these intrinsic configurations is then projected onto states with physical particle number and angular momentum to restore the symmetries broken at the mean-field level. A collective wavefunction is constructed by mixing the projected states through a variational procedure that minimizes the total energy of the system. The final nucleon configurations are then constructed with or without explicitly enforcing $\alpha$-clustering, denoted as “PGCM, C=1C=1” and “PGCM, C=0C=0”, respectively. A summary of the physically meaningful nucleon configurations obtained by the NLEFT and PGCM methods is given in Tab. 2. For more details, see the extra material provided in Ref. Giacalone and others (2025).
Figure 3: Nuclear charge density for 16O (left) and 20Ne (right) versus radius obtained with TGlauberMC for the various parameterizations discussed in the text. The respective $r_{\rm rms}$ values are given in the legend, denoted as $s$. In case of oxygen, the calculated densities are compared to the fit of the data (“Oho2”).
The nuclear charge densities for 16O obtained with TGlauberMC using the nuclear charged densities listed in Tab. 1 and the explicit nucleon configurations for NLEFT and PCGM listed in Tab. 2 in the left panel of Fig. 3, compared to the parameterization “Oho2” introduced above. As can be seen, there is a significant spread at lower radii between the different cases, which on the one hand reflect the different $\rho$ parameterizations shown in Fig. 2, and on the other hand reflects the differences between cases that implicitly or explicitly include $\alpha$-clusters (“O” and “TR_OXYGEN_V15”) exhibiting a dip at small $r$. In the right panel of Fig. 3, for completeness, meaningful profiles for 20Neon are shown, similarly revealing a large spread, with again configurations with explicit $\alpha$-clustering (“TR_NEON_V14”) exhibit a dip at small $r$ values. Not shown but mentioned in Tab. 1 are “NeTr2” and “NeTr3” because the former has a nearly 10% larger $r_{\rm rms}$ than reported from the data and the latter was fit to an unprojected version of the PCGM calculation.
V Cross section results
To compute the expected OO, NeNe and pO geometrical cross sections at 5.36 and 9.62 TeV, one needs the values for $\sigma_{\rm NN}$ at the respective nucleon–nucleon center-of-mass beam energy as input to the MCG calculation. The $\sigma_{\rm NN}$ can be obtained from the parametrization, discussed in d’Enterria and Loizides (2021),
$\sigma_{\rm NN}(\sqrt{s}/{\rm GeV})=a+b\,\ln^{n}(s/{\rm GeV}^{2})\,$, (7)
with $a=28.84\pm 0.52$, $b=0.0458\pm 0.0167$ and $n=2.374\pm 0.123$, resulting in $\sigma_{\rm NN}=68.0\pm 1.2$ mb at 5.36 TeV and $\sigma_{\rm NN}=74.7\pm 1.5$ mb at 9.62 TeV.
Figure 4: Calculated cross sections for OO collisions (left) and NeNe collisions (right) at 5.36 TeV using TGlauberMC with $d_{\rm min}=0.4$ fm, and $\omega=0.0$ (hard-sphere approximation), using the nuclear profiles and nucleon configurations as described in Sec. IV. The average as well as the average $\pm$ the standard deviation are shown as vertical lines. (The same figures including available, but non-meaningful nuclear density profiles are shown in extra Fig. 16).
The resulting cross sections for OO and NeNe collisions at 5.36 TeV are shown in Fig. 4, for $d_{\rm min}=0.4$ fm and $\omega=0$, i.e. in the hard-sphere approximation, used in previous works Loizides et al. (2018); d’Enterria and Loizides (2021). The cross sections averaged over all profiles discussed above are $\sigma_{\rm HS}=1313\pm 44$ mb and $\sigma_{\rm HS}=1678\pm 30$ mb, for OO and NeNe collisions, respectively. The above uncertainties are obtained by using the nuclear profiles as described in Sec. IV. Instead, in case of oxygen, varying the two parameters of the “Oho2” model, within the reported uncertainties would results in a relative uncertainty below 1%. The resulting cross section for pO collisions at 9.62 TeV are shown in extra Fig. 17, leading to $\sigma_{\rm HS}=466\pm 9$ mb when averaging over the various oxygen density profiles.
Recentering (discussed above) does not affect the calculated cross section, neither does changing $d_{\rm min}$ to a value lower than 0.4 fm. Increasing $d_{\rm min}$ beyond 0.4 fm, however starts to affect the calculated cross section (e.g. for $d_{\rm min}=0.8$ fm the increase is nearly 2.5%). However using $d_{\rm min}=0.8$ fm significantly distorts the charged density, which is expected given that the average minimal separation is less than 0.8 fm (evaluated using the ab-initio nucleon configurations Lonardoni et al. (2018) named “O” in TGlauberMC). Variation of $\sigma_{\rm NN}$ by $\pm 1.2$ mb, leads to a change of less than 1%.
Figure 5: Dependence of the calculated cross sections on $\omega$ for OO collisions at 5.36 TeV using TGlauberMC with the $\Gamma$-parametrization of the nucleon–nucleon overlap profile Eq. (5). Results using HIJING, PYTHIA (Monash) and TRENTO with $w=0.5$ overlap profiles are given by the thin lines.
The choice of $\omega=0$ is not really justified since particle production in pp collisions strongly depends on the nucleon–nucleon impact parameter distribution, exemplified in the development of the multi-parton picture of the PYTHIA Sjöstrand (2018) generator to describe data obtained at hadron colliders in the TeV scale. However, as discussed in Sec. III there is a large parameter space beyond the hard-sphere approximation. In particular, the calculated nucleus–nucleus cross section depends quite significantly on the nucleon–nucleon collision probability, as already discussed in Ref. Nijs and van der Schee (2022a). The dependence of the OO cross section on $\omega$ is shown in Fig. 5. The cross section is found to increase rather linearly with $\omega$ with a slope of about $161$ mb, i.e. the relative increase to the hard-sphere approximation is about $12.3\%\cdot\omega$. In addition, results from HIJING, PYTHIA and TRENTO with $w=0.5$ are indicated, which roughly span the values for $\omega$ between $0.3$ and $0.6$.
Figure 6: Dependence of the calculated cross sections on $\omega$ for PbPb, pPb, and NeNe collisions at 5.36 TeV and pO collisions at 9.62 TeV using TGlauberMC with the $\Gamma$-parametrization of the nucleon–nucleon overlap profile Eq. (5). Results using HIJING, PYTHIA (Monash) and TRENTO with $w=0.5$ overlap profiles are given by the thin lines.
Table 3: Measured and calculated cross sections for PbPb collisions at 2.76 TeV, and pPb and PbPb collisions at 5.02 TeV. The calculations using the hard-sphere approximation are from Loizides et al. (2018). The new results using the $\Gamma$-parametrization for the nucleon–nucleon overlap function with $\omega=0.3$ are obtained in this work as motivated in the text.
System Measurement Glauber (HS) Loizides et al. (2018) Glauber ($\omega=0.3$)
PbPb, $\sqrt{s_{\rm NN}}=2.76$ TeV $\sigma_{\rm PbPb}=7.7\pm 0.6$ b Abelev and others (2012) $\sigma_{\rm PbPb}^{\rm HS}=7.55\pm 0.15$ b $\sigma_{\rm PbPb}^{\rm MC}=7.70\pm 0.30$ b
PbPb, $\sqrt{s_{\rm NN}}=5.02$ TeV $\sigma_{\rm PbPb}=7.67\pm 0.25$ b Acharya and others (2024a) $\sigma_{\rm PbPb}=7.62\pm 0.15$ b $\sigma_{\rm PbPb}^{\rm MC}=7.80\pm 0.29$ b
pPb, $\sqrt{s_{\rm NN}}=5.02$ TeV $\sigma_{\rm pPb}=2.06\pm 0.03\pm 0.1$ b Khachatryan and others (2016) $\sigma_{\rm pPb}^{\rm HS}=2.09\pm 0.03$ b $\sigma_{\rm pPb}^{\rm MC}=2.13\pm 0.07$ b
$\sigma_{\rm pPb}=2.10\pm 0.07$ b Abelev and others (2014)
Table 4: Prediction cross sections for OO, NeNe, pPb and PbPb collisions at 5.36 TeV ($\sigma_{\rm NN}=68.0\pm 1.2$ mb) as well as for pO collisions at 9.62 TeV ($\sigma_{\rm NN}=74.7\pm 1.5$ mb) obtained with TGlauberMC. Given are expected values for the hard-sphere approximation and for the case of $\omega=0.3$ for the $\Gamma$-parameterization, as well as the respective relative uncertainties. Average $N_{\rm coll}$ values are obtained using $N_{\rm coll}=A\,B\,\sigma_{\rm NN}/\sigma_{\rm AB}^{\rm MC}$.
System Glauber (HS) Rel.unc. Glauber ($\omega=0.3$) Rel.unc. $\langle N_{\rm coll}\rangle$
$O$, $\sqrt{s_{\rm NN}}=5.36$ TeV $\sigma_{\rm OO}^{\rm HS}=1.31\pm 0.04$ b 3.4% $\sigma_{\rm OO}^{\rm MC}=1.36\pm 0.09$ b 6.8% 12.8
NeNe, $\sqrt{s_{\rm NN}}=5.36$ TeV $\sigma_{\rm NeNe}^{\rm HS}=1.68\pm 0.03$ b 1.8% $\sigma_{\rm NeNe}^{\rm MC}=1.73\pm 0.08$ b 4.8% 15.7
pPb, $\sqrt{s_{\rm NN}}=5.36$ TeV $\sigma_{\rm pPb}^{\rm HS}=2.10\pm 0.03$ b 2.0% $\sigma_{\rm pPb}^{\rm MC}=2.14\pm 0.07$ b 3.3% 6.6
PbPb, $\sqrt{s_{\rm NN}}=5.36$ TeV $\sigma_{\rm PbPb}^{\rm HS}=7.66\pm 0.15$ 1.5% $\sigma_{\rm PbPb}^{\rm MC}=7.80\pm 0.30$ 3.8% 340
pO, $\sqrt{s_{\rm NN}}=9.62$ TeV $\sigma_{\rm pO}^{\rm HS}=466\pm 9$ mb 1.9% $\sigma_{\rm pO}^{\rm MC}=481\pm 24$ mb 5.0% 2.3
Using an overlap function different from the hard-sphere approximation affects also PbPb and pPb collisions, as shown in Fig. 6 at 5.36 TeV, with a slope of about 495 and 139 mb, respectively. The relative increase is the same for the two systems, about 6.5%. For completeness, also NeNe collisions at 5.36 TeV, and pO collisions at 9.62 TeV are shown in Fig. 6, with slopes of 1679 and 464 mb, respectively.
To help deciding on a value for $\omega$, existing cross section data at the LHC are provided in Tab. 3 together with earlier calculations using the hard-sphere approximation Loizides et al. (2018). As can be seen, there is not a lot of room to increase the cross section beyond the hard-sphere approximation. Quite the contrary, the calculated cross sections are almost perfectly agreeing with the measurements. The recent Bayesian analysis Nijs and van der Schee (2022a) mentioned earlier, which includes a wealth of data points different from cross sections, obtained $0.4<w<0.5$ for the TRENTO model, which is similar to $\omega=0.3$ case (see Fig. 5 and Fig. 6). Indeed, the $\chi^{2}/n_{\rm dof}$ between the available data points and the $\omega$-dependent cross section obtained from the linear fits is about $0.6$ for $\omega=0.3$, and the corresponding increase is comparable in size to the systematic uncertainty quoted for the hard-sphere–based calculation. It hence is reasonable to increase the predicted cross sections by the associated relative amount (between 1.5% and 2%), and assign the same value as additional uncertainty. In this way, also the HIJING overlap profile, which qualitatively describes the centrality dependence of particle production in many collision systems Adam and others (2015); Loizides and Morsch (2017); Acharya and others (2019a); d’Enterria and Loizides (2021), is adequately covered. Using this prescription, the predicted cross sections for $\omega=0.3$ are given in Tab. 3, together with the values using the hard-sphere approximation.
VI Initial-condition related distributions
Figure 7: Number of collisions (top row) and number of participants (bottom row) versus impact parameter together for OO (left column) and NeNe (right column) collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the nuclear density profiles discussed in Sec. IV. The smaller bottom panels show the respective ratio to the average over all density profiles.
Figure 8: Participant eccentricity versus impact parameter (bottom row) for OO (left column) and NeNe (right column) collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the nuclear density profiles discussed in Sec. IV. The smaller bottom panels show the respective ratio to the average over all density profiles.
Figure 9: Ratio of participant eccentricity versus centrality for NeNe collisions to that for OO collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the ratio of the averages (standard deviation is shown as uncertainty per point), example ratio using “Ne2” and “Oho2” parameterizations as well as the ratios obtained using the nuclear densities listed in Tab. 2.
Figure 10: Ratio of participant eccentricity versus centrality for NeNe collisions to that for OO collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the ratio of the averages (standard deviation is shown as uncertainty per point), example ratio using “Ne2” and “Oho2” parameterizations as well as the ratios obtained using the nuclear densities listed in Tab. 2. Gaussian smearing with $p=1$ (left) and $p=0$ (right panel) for the nuclear overlap is used as explained in the text.
In the following, the number of participants ($N_{\rm part}$) typically needed to compare measured quantities across systems and number of collisions ($N_{\rm coll}$) typically used to quantify the nuclear modification factor are discussed. Average $N_{\rm coll}$ can be obtained using $N_{\rm coll}=A\,B\,\sigma_{\rm NN}/\sigma_{\rm AB}^{\rm MC}$ leading to $12.8\pm 0.8$, $15.8\pm 1.1$, $2.49\pm 0.12$ for OO and NeNe at 5.36 TeV and pO at 9.62 TeV (see Tab. 4). In case of pO (or pPb) collisions, it follows that average $N_{\rm part}=\langle N_{\rm coll}\rangle+1$.
Figure 7 shows the $N_{\rm coll}$ and $N_{\rm part}$ distributions versus the impact parameter $b$ for OO and NeNe collisions at 5.36 TeV computed with TGlauberMC ($d_{\rm min}=0.4$fm, $\omega=0.3$) using the nuclear density profiles discussed in Sec. IV. In more central collisions (smaller impact parameter) the $N_{\rm coll}$ values differ with respect to the average by 5% for oxygen and up 10% for neon. In peripheral collisions (larger impact parameter) the differences increase to 10% and 15%, respectively, and a grouping into different types of underlying nuclear density profiles (Wood-Saxon, NLEFT, PGCM) becomes apparent. For $N_{\rm part}$ the differences are smaller, on the level of few percent in central collisions, reaching 5% for oxygen and 10% for neon in peripheral collisions. In the actual measurements, the observables will be measured in intervals of multiplicity and/or energy, which for large nuclei are well correlated with centrality. The intervals then will be mapped to $N_{\rm part}$ (and $N_{\rm coll}$) with different procedures summarized in Ref. Miller et al. (2007); d’Enterria and Loizides (2021), which rely on MCG calculations. The process requires the data and calculations to be matched at the so-called anchor point. In PbPb collisions the anchor point can be determined rather precisely to the 1% level, and up to 90% peripheral events Abelev and others (2013). For OO (and NeNe) collisions, this may not be achievable, due to the much larger fluctuations and the unknowns of the underlying nuclear density profile.
Figure 8 shows the participant eccentricity versus the impact parameter $b$ for OO and NeNe collisions at 5.36 TeV. Both exhibit an increasing trend from central to peripheral collisions reflecting the increasing elliptical shape of the overlap region. Due the significant differences in the nuclear profiles, there are differences of up to 15% relative to the average over all profiles, in particular in central collisions.
Figure 9 shows the ratio of participant eccentricity for NeNe collisions to that for OO collisions at 5.36 TeV versus centrality (estimated from the respective impact parameter distributions) for the ratio of the averages (including the standard deviation as uncertainty) as well as direct ratios from the nuclear densities listed in Tab. 2. As can be seen, the average ratio rises with increasing impact parameter, reaching about 1.05 in central collision. On the other hand, the NLEFT and PGCM profiles reach values significantly larger values up to 1.10–1.15. Assuming an hydrodynamical expansion of a nearly inviscid fluid, $\varepsilon_{2}\propto v_{2}$ holds to first order Noronha-Hostler et al. (2016), the ratio of $v_{2}$ between oxygen and neon is predicted to be of similar size. Indeed, a detailed calculation Giacalone and others (2025) expects the ratio to be $1.15$ to within $\pm 3\%$ for most central collisions, since most theoretical uncertainties cancel out in the $v_{2}$ ratio.
Figure 11: Initial-state $\varepsilon_{3}$ (top) and $\varepsilon_{4}$ (bottom) versus impact parameter for OO (left column) and NeNe (right column) collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the nuclear density profiles discussed in Sec. IV. The smaller bottom panels show the respective ratio to the average over all density profiles.
By construction, the participant eccentricity approaches unity in the most peripheral collisions when it is calculated directly from the participant positions, which in the limiting case may consist of only two participants. This behavior was first investigated in Ref. Alver and others (2008), which demonstrated for AuAu collisions that applying (Gaussian) smearing to the participant positions primarily affects the participant eccentricity in peripheral events (up to approximately $N_{\rm part}\sim 30$). For small systems, however, this range effectively covers the entire domain of interest, and smearing therefore has a substantial impact. To study whether smearing has an effect on the predicted ratio $\varepsilon_{2}^{\rm NeNe}/\varepsilon_{2}^{\rm OO}$, one commonly relies on the generalized mean for the nuclear overlap function defined as $\left(\frac{(T_{A}^{p}+T_{B}^{p})}{2}\right)^{\frac{1}{p}}$ Moreland et al. (2015) which for $p=1$ reduces to the arithmetic mean and for $p=0$ to the geometric mean (also known as TRENTO initial conditions). As shown in Fig. 10, in both cases the ratio approaches similar large values towards most central collisions as without Gaussian smearing. However, the trend with decreasing centrality differs, in particular for the TRENTO ($p=0$) condition, which increases with decreasing centrality.
For completeness, the distributions for $\varepsilon_{3}$ and $\varepsilon_{4}$ versus impact parameter for OO and NeNe collisions at 5.36 TeV are shown in Fig. 11 using the nuclear density profiles listed in Tab. 2. Both exhibit an increasing trend from central to peripheral collisions as known from larger nuclei collision systems. Variations of up to 10% relative to the average are observed as a result of the differences between the nuclear density profiles. Figure 12 shows the ratios of $\varepsilon_{3}$ and $\varepsilon_{4}$ for NeNe collisions to that for OO collisions at 5.36 TeV for the ratio of the averages (including the standard deviation as uncertainty) as well as direct ratios from the nuclear densities listed in Tab. 2. The ratio of $\varepsilon_{3}$ between NeNe and OO exhibits a modest decrease with decreasing centrality, turning into a strong rise for most peripheral collisions. The ratio of $\varepsilon_{4}$ between NeNe and OO changes only little with centrality, most notably with a decreasing trend with decreasing centrality for the NLEFT calculation by about 5 percent.
Figure 12: Ratio of $\epsilon_{3}$ (left panel) and $\epsilon_{4}$ (right panel) versus centrality for NeNe collisions to that for OO collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the ratio of the averages (standard deviation is shown as uncertainty per point), example ratio using “Ne2” and “Oho2” parameterizations as well as the ratios obtained using the nuclear densities listed in Tab. 2.
VII Particle production
The centrality dependence of charged-particle production in heavy-ion collisions was found to scale with the number of participants Alver and others (2011). In particular at LHC energies Loizides (2011), inclusion of a small contribution from the number of collisions improves the description to account for the increase of multiplicity in more central collisions from mini-jets, as ${\rm d}N/{\rm d}\eta=\alpha N_{\rm part}+\beta N_{\rm coll}$. Typically, ${\rm d}N/{\rm d}\eta$ is normalized per participant pair, $N_{\rm part}/2$.
When this formalism is applied across different collision energies and system sizes, employing sub-nucleon degrees of freedom by decomposing each nucleon into three or more effective quark-level interaction centers—removes the need to reference the number of binary nucleon–nucleon collisions, as approximate scaling arises from the generalized number of participant quark centers (see Loizides (2016b) and references therein).
Figure 13: Average $N_{\rm mpi}/N_{\rm part}$ (left panel) and $R\equiv\left\langle N_{\rm mpi}\right\rangle/(\left\langle N_{\rm coll}\right\rangle\,\left\langle N_{\rm mpi}^{\rm pp}\right\rangle)$ (right panel) versus $N_{\rm part}$ for OO, NeNe, XeXe and PbPb collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$.
Here, I take a different approach, and introduce the number of multiple hard interactions as implemented in HIJING Wang and Gyulassy (1991) and applied recently in Loizides and Morsch (2017); Jonas and Loizides (2021). For a given nucleon–nucleon collision the number of hard scatterings at $b_{\rm NN}$ is distributed according to a Poisson distribution with the average given by $\langle N^{\rm mpi}_{\rm NN}\rangle=\sigma_{\rm NN}^{\rm hard}\,T_{\rm NN}(b_{\rm NN})$, where $\sigma_{\rm NN}^{\rm hard}$ is the energy-dependent pQCD cross-section for $2\to 2$ parton scatterings ($\approx 190$ mb at 5.36 TeV). The total number of multiple interactions $N_{\rm mpi}$ is then the sum of the multiple interactions over all nucleon–nucleon collision in the event, and on average equal to $N_{\rm coll}$ times the average $N_{\rm mpi}$ in pp collisions at the same energy. Since the $N_{\rm mpi}$ and event multiplicity are strongly correlated Loizides and Morsch (2021), ordering events according to $N_{\rm mpi}$ can be used at a proxy for event categorization usually done to group events into centrality intervals. In this case, the relation on the average between $N_{\rm mpi}$ and $N_{\rm coll}$ does no longer need to hold. The left panel of Fig. 13 shows the resulting $\left\langle N_{\rm mpi}\right\rangle$ distributions normalized by $\left\langle N_{\rm part}\right\rangle$ in centrality-binned intervals versus $N_{\rm part}$ for OO, XeXe and PbPb collisions at 5.36 TeV. After a steep rise at very low $N_{\rm part}$, the $N_{\rm mpi}$ in the different systems rise approximately linearly with $N_{\rm part}$, until they eventually increase faster than linear with $N_{\rm part}$, most evidently seen for OO and XeXe collisions. In the right panel, $R\equiv\left\langle N_{\rm mpi}\right\rangle/(\left\langle N_{\rm coll}\right\rangle\,N_{\rm mpi}^{\rm pp})$ versus $N_{\rm part}$ is shown, which exhibits the effect of the geometry and event selection bias on the nuclear modification factor, even in absence of any nuclear effect Loizides and Morsch (2017). Indeed, the ratio significantly deviates from one in peripheral collisions, where the ratio is smaller up to about 50%, and while it is enhanced by up to 10% in the most central OO collisions. In reality, effects from fragmentation, momentum conservation, play a role, which are expected to result in a $p_{\rm T}$-dependent selection bias Park et al. (2025).
Figure 14: Per-participant charged particle pseudo-rapidity density $2/\left\langle N_{\rm part}\right\rangle\,{\rm d}N/{\rm d}\eta$ versus $\left\langle N_{\rm part}\right\rangle$ at mid-rapidity in XeXe collisions at 5.44 TeV (left) and in PbPb collisions at 5.36 TeV (right) obtained by ALICE and CMS Acharya and others (2019b); Sirunyan and others (2019); Hayrapetyan and others (2025d); Acharya and others (2025), as well as two component fits to each dataset (see text). Data Acharya and others (2023) for pp collisions normalized to non-single diffractive and inelastic yields are also shown.
Per-participant charged particle pseudo-rapidity density $2/\left\langle N_{\rm part}\right\rangle\,{\rm d}N/{\rm d}\eta$ versus $\left\langle N_{\rm part}\right\rangle$ at mid-rapidity in XeXe collisions at 5.44 TeV and in PbPb collisions at 5.36 TeV obtained by ALICE and CMS Acharya and others (2019b); Sirunyan and others (2019); Hayrapetyan and others (2025d); Acharya and others (2025) are shown in Fig. 14. In case of XeXe collisions the ALICE and CMS date are not consistent for peripheral collisions. Data for pp collisions normalized to non-single diffractive and inelastic yields are also shown (scaled to 5.36 TeV by less than one percent as given by respective power laws as a function of $\sqrt{s}$Acharya and others (2023).
Figure 15: Per-participant charged particle pseudo-rapidity density $2/\left\langle N_{\rm part}\right\rangle\,{\rm d}N/{\rm d}\eta$ versus $\left\langle N_{\rm part}\right\rangle$ at mid-rapidity in XeXe collisions at 5.44 TeV and in PbPb collisions at 5.36 TeV obtained by ALICE and CMS Acharya and others (2019b); Sirunyan and others (2019); Hayrapetyan and others (2025d); Acharya and others (2025), as well as data for pp collisions normalized to non-single diffractive and inelastic yields Acharya and others (2023). The figures zooms into the low $N_{\rm part}$ region to show highlight predictions for OO and NeNe collisions at 5.36  TeV obtained by average of the fits obtained in Fig. 14 together with the respective two-component model fits (see text).
Despite the apparent opposite trend in the PbPb case the dataset are consistent considering the large uncertainties in peripheral collisions. Since the data are essentially at the same collision energy as that for the planned OO collisions, they can be used to predict the ${\rm d}N/{\rm d}\eta$ for OO collisions, by fitting a two-component model $2N(f+(1-f)\,{\left\langle N_{\rm mpi}\right\rangle}/{\left\langle N_{\rm part}\right\rangle})$. The fits are found to describe the data reasonably well, however with different values for $N$ between XeXe and PbPb collisions, in particular for ALICE. Hence, for the predicted result, the central value between the ratios is used with half of the spread as uncertainty assigned to essentially cover all cases. The same procedure was also applied to give the expected mid-rapidity ${\rm d}N/{\rm d}\eta$ for NeNe collisions. The expectations for OO and NeNe collisions at 5.36 TeV are shown in Fig. 15 together with the existing data.
Table 5: Centrality dependence of $N_{\rm part}$, $N_{\rm coll}$, $N_{\rm mpi}$, $R\equiv\left\langle N_{\rm mpi}\right\rangle/(\left\langle N_{\rm coll}\right\rangle\,\left\langle N_{\rm mpi}^{\rm pp}\right\rangle$ and $\frac{2}{N_{\rm part}}{\rm d}N/{\rm d}\eta/N_{\rm part}$ in OO and NeNe collisions at 5.36 TeV. The Glauber values are given without uncertainties.
$O$ NeNe
Centrality $\langle N_{\rm part}\rangle$ $\langle N_{\rm coll}\rangle$ $\langle N_{\rm mpi}\rangle$ $R$ $\frac{2}{N_{\rm part}}{\rm d}N/{\rm d}\eta$ $\langle N_{\rm part}\rangle$ $\langle N_{\rm coll}\rangle$ $\langle N_{\rm mpi}\rangle$ $R$ $\frac{2}{N_{\rm part}}{\rm d}N/{\rm d}\eta$
[…]
0–100% 10.8 12.8 37.6 1 $6.02\pm 0.21$ 12.7 15.7 46.0 1 $6.06\pm 0.21$
A table with all relevant Glauber related values and projected multiplicities is given in Tab. 5. The Glauber values are given without uncertainties. However, it should be noted that the associated uncertainties usually estimated from data require careful evaluation, and are expected to be non-negligible.
VIII Summary
An updated version (v3.3) of the TGlauberMC model is presented, which includes several different probability distributions for the nucleon–nucleon interaction probability (Fig. 1), and a huge variety of nuclear density profiles (Tab. 1 and 2), in particular for oxygen and neon (Fig. 2 and 3). The model is used to predict properties of OO, NeNe, as well as pPb and PbPb collisions at 5.36 TeV, and pO collisions at 9.62 TeV. The expected geometrical cross sections are averaged over all provided profiles (Fig. 4), and considered to include a contribution from the transverse nucleon–nucleon overlap (Fig. 5 and  6), and presented in Tab. 4. For central collisions, the distributions of $N_{\rm coll}$ and $N_{\rm part}$ exhibit a spread of 5–10% and about 2%, respectively (Fig. 7), and up to 20% for the participant eccentricity (Fig. 8). The ratio of $\varepsilon_{2}$ in NeNe to OO is found to reach up to $1.15\pm 0.05$ (Fig. 9), rather independently whether Gaussian smearing (Fig. 10) using generalized TRENTO conditions is used or not. Rising trends for $\varepsilon_{3}$ and $\varepsilon_{4}$ with decreasing centrality (Fig. 11), while first decreasing then rising trends with decreasing centrality of the respective ratios between NeNe and OO collisions (Fig. 12) are found. Using $N_{\rm mpi}$ instead of $N_{\rm coll}$ (Fig. 13) allows to predict general trends of the OO data (Fig. 14) and use existing data at 5.36 TeV to predict the multiplicity in OO and NeNe collisions (Fig. 15). The source code for TGlauberMC (v3.3) is available at HepForge 92, and the author welcomes suggestions how to improve it.
During the peer-review of the manuscript results from OO collisions at $\sqrt{s_{\rm NN}}=200$ GeV, taken in June, 2021 at RHICAboona and others (2025), and from OO and NeNe collisions at $5.36$ TeV at LHC I. J. Abualrob et al. (2025); G. Aad et al. (2025); A. Hayrapetyan et al. (2025b); 69; A. Hayrapetyan et al. (2025a); 34 were submitted. The availability of the OO data at both RHIC and LHC, opens an important opportunity to explore both particle production and collective flow across a wide energy range — and TGlauberMC (v3.3) should prove a valuable tool for interpreting those measurements.
Acknowledgements.
The author would like to thank G. Nijs and W. van der Schee for fruitful discussions related to nuclei configurations provided by Trajectum, G. Giacalone for clarification related to the 3pF fit of the NNLOsat oxygen profile, V. Soma for providing the exact data points of the NNLOsat calculation Somà et al. (2020), B. Kardan for contributing changes in TGlauberMC v3.3 useful for Hades/CBM collision energies, as well as F. Jonas and N. Strangmann for taking the time to read and comment on a draft of this article. The author acknowledges financial support by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under contract number DE-SC0005131.
Appendix A TGlauberMC v3.3 user’s guide
The main changes related to the previous version (3.2) is the interface to any nucleus configuration provided by Trajectum Nijs and van der Schee (2022b) via the new TrNucGen 93 library (see App. B), the support to read nucleus configurations from text files, and a larger list pre-defined nucleus configurations. Further, the possibility to use between overlap functions using TRENTO, HIJING and PYTHIA parameterizations, in addition to the $\Gamma$-distribution that already was present since v3.0. As default, the code continues to use the hard-sphere approximation ($\omega=0$, see Sec. III). Following the discussion in Sec. VII, the total number of multiple interactions are computed from the nucleon–nucleon overlap function. The $\sigma_{\rm NN}$ parameterization for RHIC and LHC energies is provided, as well as dedicated parameterizations Bystricky et al. (1987) for neutron–neutron (proton–proton) and neutron–proton interactions relevant at Hades/CBM energies.
The user’s guide has been incrementally updated from the first release of the code in 2008 Alver et al. (2008), the improved version in 2014 Loizides et al. (2015), and the latest in 2017 Loizides et al. (2018). As before, the source code, which relies on the ROOT Brun and Rademakers (1997) framework (version 6.x), can be obtained at the TGlauberMC page on HepForge 92. For version 3.3, it also provides a “doxygen” documentation, which provides many details. In the following, I only describe the improved functionality relative to the last update Loizides et al. (2018).
All functionality is implemented in the macro runglauber_vX.Y.C, where version “X.Y==3.3” described here. It is best to use the provided “rootlogon.C” macro in the same directory which will setup the code automatically. Unlike before, additional input files for nucleus configurations are now located in the subdirectory “dat”. Three classes, TGlauNucleon, TGlauNucleus and TGlauberMC and five functions (macros) runAndSaveNtuple(), runAndSaveNucleons(), runAndSmearNtuple(), runAndOutputLemonTree() and runAndCalcDens() are defined in the provided macro as in earlier versions.
As before, the most convenient starting point is the provided runAndSaveNtuple() which will generate the requested number of MCG events for given nucleus types, $\sigma_{\rm NN}$ as well as most notably the $d_{\rm min}$ and $\omega$ values, storing the event-by-event computed quantities in a ROOT tree. The only difference to earlier versions is that if the provided value for $\sigma_{\rm NN}$ is negative, it will be interpreted as beam energy to compute $\sigma_{\rm NN}$ and $\sigma_{\rm NN}^{\rm hard}$. The updated list of supported nuclei can be found in the TGlauNucleus::Lookup function in the code. Nuclei names that start with “TR” are taken from TrNucGen, see App. B. For this to work, the TrNucGen library must be available in a subdirectory called “trnucgen” (or a symbolic link must be present that points to it). Nuclei names that start with “input” are read from a text file with the following structure:
# Nucleus generator version: 1.0.0
# Nucleus type: TR_CARBON_V1
# Nucleus id: 7
# Nucleus name: C1
# Nucleus description: carbon12(v1)
# Number of nucleons: 12
# Number of protons: 6
# Number of events: 10000
# Per line all nuclei, in the format x1, y1, z1, c1, x2, y2, z2, c2, ...
where the position ($x$, $y$, $z$) and charge $c$ are listed per line for a nucleus configuration. The format of the file is also explained in the example provided with TrNucGen 93).
The following set of functions controls additional behavior of the TGlauNucleus class: GetDens returns the density normalized to $Z$ or to $A$. GetSqrtMeanR2() returns the $r_{\rm rms}$ computed from the density distribution. CalcMinDist() calculates the smallest difference between any two nucleons in the nucleus. CalcRmsRadius() calculates the $r_{\rm rms}$ computed from the nucleon positions assigned to the nucleus. GetWeight and SetWeight() returns or sets the weight for the particular nucleon configuration (only used by nuclei constructed with TrNucGen). SetBeta(Double_t b2, Double_t b3, Double_t b4, Double_t g) allows to set an almost deformation parameters $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ and $\gamma$ that together with “box” Monte Carlo sampling allows the generation for deformed nuclei.
The following set of member functions controls additional behavior of the TGlauberMC class: SetOmega($w$) to provide the nucleon–nucleon profile to use. The input value can be $0<w<2$ for the $\Gamma$-distribution, $7$ for HIJING, $8$ for PYTHIA and $9<w<11$, which would translate into $0<\delta<2$ for TRENTO. SetSigmaHard to set the hard scattering cross section (in case you do not want to rely on the precomputed value. SetShadowing to control whether or not shadowing corrections are computed.
In addition to quantities described the previous user guides, the following quantities are stored in the ROOT tree:
  • Nmpi: Number of MPI (HIJING model)
  • NpartAn: Number of wounded (participating) neutrons in Nucleus A
  • NpartBn: Number of wounded (participating) neutrons in Nucleus B
  • Npart0n: Number of singly-wounded (participating) neutrons
  • SpecA: Spectator neutrons in nucleus A
  • SpecB: Spectator neutrons in nucleus B
  • Weight: Weight of event (needed for e-by-e weighting)
Additional global functions that are provided are
  • getNNProf() and getNNProfDist for the $\Gamma$-based collision probability
  • getNNHijing() and getNNHijingDist() for the HIJING collision probability
  • getNNPythia() and getNNPythiaDist() for the PYTHIA collision probability
  • getNNTrento() and getNNTrentoDist() for the TRENTO collision probability
  • getSigmaNNvsEnergy() to get $\sigma_{\rm NN}$ as function of collision energy
  • getSigmaNN() to return the $\sigma_{\rm NN}$ for a given energy
  • getSigmaNPvsEnergy_Bystricky() and getSigmaPPvsEnergy_Bystricky() to get $\sigma_{\rm NN}$ and $\sigma_{\rm NP}$ as function of energy (useful mainly below 10 GeV)
  • getSigmaNP_Bystricky() and getSigmaPP_Bystricky() to get $\sigma_{\rm NN}$ and $\sigma_{\rm NP}$ (useful mainly below 10 GeV)
  • getSigmaHardvsEnergy() returns $\sigma_{\rm NN}^{\rm hard}$ as a function of energy
  • getSigmaHard() returns $\sigma_{\rm NN}^{\rm hard}$ at the given energy
These also exists as corresponding static functions of the TGlauberMC class. Additional static functions are ReadNtuple to read the produced Ntuple, GetCentralityDist to produce a centrality distribution based on impact parameter and AddCentralityBranch to add the centrality estimator to the produced tree.
The runAndSmearNtuple() function has been improved to include generalized overlap function defined as $\left(\frac{(T_{A}^{p}+T_{B}^{p})}{2}\right)^{\frac{1}{p}}$ Moreland et al. (2015), which for $p=1$ reduces to the arithmetic mean and for $p=0$ to the geometric mean (also known as TRENTO initial conditions). The function produces a sample of $n$ events with a generalized reduced-thickness (or $p$-parameter) formulation of the overlap. The parameter $p$ controls the functional form of the participant thickness; unless stated otherwise the default value $p=1$ is used, while $p=0$ corresponds to the TRENTO prescription. The nuclei involved in the collision are specified by the strings sysA and sysB, which are interpreted through the TGlauNucleus::Lookup tables. The inelastic nucleon–nucleon cross section is set by $\sigma_{NN}$ (in units of mb), for which a default value of $68\,\mathrm{mb}$ is used. A minimum nucleon–nucleon separation $d_{\min}$ is enforced during nucleon placement, with $d_{\min}=0.4\,\mathrm{fm}$ in our baseline configuration. The model allows for a tunable nucleon profile characterized by the parameter $\omega$. A value $\omega=0$ corresponds to a hard-sphere profile, $0\leq\omega\leq 2$ yields a Gamma-distributed profile, and $\omega=7$ and $8$ reproduce the effective profiles used in HIJING and PYTHIA, respectively. Larger values ($\omega=9$$11$) correspond to the generalized TRENTO implementations. Additional profile-shape parameters include $w$, for which the TRENTO-motivated value $w=0.5$ is used, as well as $k$, the shape parameter of the Gamma distribution, with default value $k=1.4$. Events are sampled within an impact-parameter range $b_{\min}\leq b\leq b_{\max}$, where $b_{\min}=0$ and $b_{\max}=20\,\mathrm{fm}$ are used unless otherwise stated. The resulting events are written to an output ntuple specified by fname; if no file name is supplied, a default name is automatically generated.
Appendix B TrNucGen
The Trajectum Nucleus Generator (TrNucGen) software library 93 provides all nuclei configurations available in Trajectum v2.1 Nijs and van der Schee (2022b). It was written in particular to use all the configurations for Oxygen-16 and Neon-20 nuclei Giacalone and others (2025) explored in this paper. The library is defines a convenient interface to the relevant code extracted from Trajectum, avoiding having to compile the latter with all the required dependencies. The full list of available oxygen and neon profiles from Trajectum is given in Tab. 6. The list of all nuclei configurations can be found in the doxygen documentation of the code. The installation instructions as well as the source code including a standalone example are publicly available 93. The library can be also very useful for any heavy-ion model or simulation that would like to explore nuclei configurations beyond those based on common parameterizations (i.e. 3pF).
Table 6: Full list of oxygen-16 and neon-20 NLEFT and PGCM calculations taken from Trajectum Giacalone and others (2025).
Profile Description
TR_OXYGEN_V1 Variational Monte Carlo version Lim et al. (2019) (same as “O” in TGlauberMC v2.7 and higher)
TR_OXYGEN_V2 NLEFT pinhole, $\pm$ weights
TR_OXYGEN_V3 3pF from Summerfield et al. (2021) (same as “Opar” in TGlauberMC v2.7 and higher)
TR_OXYGEN_V4 3pF fit to NNLOSat Somà et al. (2020) (same as “Opar2” in TGlauberMC v3.3)
TR_OXYGEN_V5 Ab-initio unprojected PGCM.
TR_OXYGEN_V6 NLEFT version with only positive weights
TR_OXYGEN_V7 Ab-initio unprojected PGCM, enforcing clustering
TR_OXYGEN_V8 Ab-initio unprojected PGCM, chopping off artifact regions
TR_OXYGEN_V9 Ab-initio unprojected PGCM, enforcing clustering and chopping off artifact regions.
TR_OXYGEN_V10 PGCM, constraints imposed on unprojected states, no explicit clustering
TR_OXYGEN_V11 PGCM, constraints imposed on unprojected states, with explicit clustering
TR_OXYGEN_V12 NLEFT pinhole, $\pm$ weights, with periodicity ambiguities resolved
TR_OXYGEN_V13 NLEFT version with only positive weights, with periodicity ambiguities resolved
TR_OXYGEN_V14 PGCM, constraints imposed on projected states, no explicit clustering
TR_OXYGEN_V15aaa PGCM, constraints imposed on projected states, with explicit clustering
TR_NEON_V1 Ab-initio unprojected PGCM
TR_NEON_V2 3pF plus deformation (same as “NeTr2” in TGlauberMC v3.3
TR_NEON_V3 Fit to TR_NEON_V1 (same as “NeTr3” in TGlauberMC v3.3
TR_NEON_V4 NLEFT pinhole, $\pm$ weights
TR_NEON_V5 NLEFT version with only positive weights
TR_NEON_V6 Ab-initio unprojected PGCM, enforcing clustering
TR_NEON_V7 Ab-initio unprojected PGCM, chopping off artifact regions.
TR_NEON_V8 Ab-initio unprojected PGCM, enforcing clustering and chopping off artifact regions
TR_NEON_V9 PGCM, constraints imposed on unprojected states, no explicit clustering
TR_NEON_V10 PGCM, constraints imposed on unprojected states, with explicit clustering
TR_NEON_V11 NLEFT pinhole, $\pm$ weights, with periodicity ambiguities resolved
TR_NEON_V12 NLEFT with only positive weigths, with periodicity ambiguities resolved
TR_NEON_V13 PGCM, constraints imposed on projected states, no explicit clustering
TR_NEON_V14 PGCM, constraints imposed on projected states, with explicit clustering
Appendix C Additional figures
This section shows additional figures only available in the arXiv version of this text, and mainly for completeness.
Figure 16: Calculated cross sections for OO and NeNe collisions at 5.36 TeV using TGlauberMC with $d_{\rm min}=0.4$ fm and $\omega=0$ (hard-sphere) approximation. The results shown in gray were obtained with non-meaningful profiles, and are not included in the determination of the average and standard deviation.
Figure 17: Calculated cross sections for pO collisions at 9.62 TeV using TGlauberMC with $d_{\rm min}=0.4$ fm and $\omega=0$ (hard-sphere) approximation. The right panel shows in gray also additional results obtained with non-meaningful profiles for oxygen, and are not included in the determination of the average and standard deviation.